OpenStudy (anonymous):

which algebraic expression is equivalent to the expression bellow? (the numbers next to the / sign means they're joined together as a fraction) 2/5(8x + 5 + 3x) A)22/5x + 2 B) 48/5x + 2 C) 32/5x

OpenStudy (texaschic101):

(2/5)(8x + 5 + 3x) = (2/5)(11x + 5) = (22/5)x + 10/5 = 22/5x + 2

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

nope that's all thanks so much :) oh do you know how to give medals... i have to give you one... you help me i help you

OpenStudy (anonymous):

oh nevermind i just gave you one :)

OpenStudy (texaschic101):

to give medals, you click on best response by whoever you want to give the medal to. Glad I could help :)

OpenStudy (anonymous):

thanks bye:)